# Planck dust polarization power spectra are consistent with strongly supersonic turbulence

### v1

@@ 3.2 Projected Intensity and Polarization Power Spectra
The $T$ spectrum, $C_{k}^{TT}$, can be seen in the first panel of Figures 6 and its slope, $\alpha_{TT}$, in the first panel of Figure 7. The $T$ spectrum has a slope that is nearly identical to the $\rho$ spectrum (though they appear different due to the difference in compensation). This is expected as in this model, $T$ is simply the projection of $\rho$. By the slice-projection theorem, given a quantity $q(x,y,z)$, its projection $Q(x,y)$, and their Fourier transforms $\hat{q}(k_{x},k_{y},k_{z})$ and $\hat{Q}(k_{x},k_{y})$ are related as
@@
Refer to captionRefer to caption
Figure 9: (Top) Correlation ratio $r_{xy}$ for TE, TB, and EB spectra. Vertical bars show the analysis window. (Bottom) Average $r_{XY}$ in the analysis window versus Sonic Mach number KMH: label missing, $\mathcal{M_{S}}$, compared to Planck-measured values. In the top plots, we use dotted lines for $0.70.4<\mathcal{M_{\rm{A}}}<$ 0.7 (smallest point size at bottom), dashed lines for $1.20.7<\mathcal{M_{\rm{A}}}<$ 1.2, and solid lines for $2.21.2<\mathcal{M_{\rm{A}}}<$ 2.2 (largest point size at bottom). See text for cautions about interpreting the $\langle r_{TB}\rangle$ and $\langle r_{TE}\rangle$.
@@ 3.3 Power Amplitude Ratios
The ratio of $B$ power to $E$ power is interest because the observed value of $A_{\rm{BB}}/A_{\rm{EE}}=0.53$ was unexpected. We can also examine the ratio of of $E$ and $B$ relative to total power, $A_{\rm{EE}}/A_{\rm{TT}}$ and $A_{\rm{BB}}/A_{\rm{TT}}$, though our model lacks a detailed treatment of the dust polarization fraction, which implies that these ratios cannot be directly compared to observed values, though the trend with $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ can be measured. The polarization fraction cancels out in the $A_{\rm{BB}}/A_{\rm{EE}}$ ratio, so this can be directly compared to the Planck value.
Figure 8 shows the ratio of fit amplitudes, $A_{\rm{EE}}/A_{\rm{TT}}$, $A_{\rm{BB}}/A_{\rm{TT}}$, and $A_{\rm{BB}}/A_{\rm{EE}}$ versus the sonic Mach number in each of the three panels. The third panel also shows the observed value as a grey line. The runs with slower velocities show the most $E$ modes, as the low velocity causes the field to have larger impact on the morphology, and more filamentary structures that align magnetic field and density are observed. This can be seen in the projections in Figure 1. The runs with the slower velocities show the largest $B$ modes as well, when the Alvèn Mach number is large and the magnetic fields are weak. But when then the magnetic fields are strong, the slower velocities show the weakest B modes. For higher velocity Mach numbers ($\mathcal{M_{S}}\geq 4$), the amplitude ratios depend little on either the fluid velocity or the magnetic field. For $\mathcal{M_{S}}>4$, the ratio tends toward $A^{BB}/A^{EE}=0.55\pm 0.07$, near to the observed ratio of Planck. Thus it may be that that the ratio of $B$ to $E$ observed by Planck is a natural consequence of compressive turbulence, and that $A_{\rm{BB}}/A_{\rm{EE}}=0.53$ simply because the flow is hypersonic and magnetized, which naturally gives this value.
@@ 3.4 Cross-correlations
for all pairs of $T$, $E$, and $B$. The top row shows the spectra, with the fitting window denoted by vertical light-grey lines. The bottom row shows $\langle r^{\rm{XY}}\rangle$, averaged over the fit window and all frames, as a function of Mach number. Error bars are found by first averaging $r_{k}^{XY}$ over $k$ for each frame within the fit window, then taking the variance over frames. Shown in horizontal grey lines are the observed values of $r^{\rm{TE}}=0.355$ and $r^{\rm{TB}}=0.055$ (Planck Collaboration et al., 2020).
For the two even-parity modes, $T$ and $E$, we find significant positive correlation in all cases but one. As the velocity decreaess and magnetic field increaes, $r^{\rm{TE}}$ increases well above the value observed by Planck. The general trend is of increasing correlation with increasing field strength. For large Mach numbers, correlations are more modest and mostly consistent with the observed Planck value of $r^{\rm{TE}}=0.35$. For large $\mathcal{M_{S}}$, the effect of the magnetic field is not as pronounced as it is for lower $\mathcal{M_{S}}$. Earlier results, e.g. for $\alpha_{EE}$ and the ratio of $E$ to $B$, show that the Planck data are consistent with $\mathcal{M_{S}}>4$, and this is also compatible with what we see for $\langle r^{\rm{TE}}\rangle$. We do observe a scale dependence in $r^{\rm{TE}}$ (rising toward small scales) that is not seen by Planck, which may be a sign that our simplified simulations are missing a key element of the ISM.
How do we draw conclusions from the $\langle r^{\rm{TE}}\rangle$ values, compared to the Planck value of 0.35? The Planck correlation looks most compatible with slow velocities with moderate magnetic fields or fast velocities with moderate-to-strong magnetic fields. Earlier we saw that the $B/E$ power ratio prefers fast velocities with $\mathcal{M_{S}}>4$, so the latter case seems to fit the bill. Other combinations do not work: in our simulations, low $\mathcal{M_{S}}$ and low $\mathcal{M_{\rm{A}}}$ (low velocity and strong magnetic field) would lead to a stronger correlation than what Planck sees. Low $\mathcal{M_{S}}$ and high $\mathcal{M_{\rm{A}}}$ (low velocity and weak magnetic field) leads to too small of a correlation, or a slight anticorrelation.
@@
These tendencies toward positive parity-violating correlation values at high fluid velocity are somewhat puzzling because all of the MHD physics we include respects parity. The simulations start with uniform density and are driven with a non-helical acceleration pattern, and the base solver is an unsplit solver with no inherent asymmetry. A larger, more systematic, and more resolved ensemble of simulations will be necessary to determine if we have inadvertently inserted some parity-violating effect, if this is simply a statistical fluctuation, or if there is a slight tendency for MHD simulations to produce positive parity violations.
@@ 3.5 Linear fits and importance of parameters
where $q$ stands for density, velocity, magnetic field, $T$, $E$, and $B$. These fit coefficients are found in Table 1. Noting that $b=\frac{\partial q}{\partial M_{s}}$ and $c=\frac{\partial q}{\partial M_{A}}$, we see that $b$ and $c$ give the relative importance of sonic and Alfvén Mach numbers on each quantity. The third column of Table 1 gives the ratio of $c/b$, which denotes the relative impact of the two. Density, $\rho$, depends only on sonic Mach number $(b_{\rho}=0)$ while the velocity spectrum is more influenced by $\mathcal{M_{\rm{A}}}$ ($c_{\rho}/b_{\rho}=6.5$). Sonic Mach number determines $\alpha_{TT}$, while Alfvén Mach number determines $\alpha_{BB}$.
@@ 3.6 Parallel versus perpendicular projections
Figure 11 shows $\alpha_{EE}$ and $\alpha_{BB}$ vs $\mathcal{M_{S}}$ for parallel projections (top row) and perpendicular projections (bottom row, same as Figure 7, reproduced for ease of comparison). For the weakly magnetized cases (large points) the behavior is comparable between the two directions, as expected. For the more strongly magnetized case, increasing magnetic field has the opposite effect on the slope between the two directions. For the $\hat{x}$ projection, increasing magnetic field increases $\alpha_{EE}$ and $\alpha_{BB}$ slightly. For the parallel direction, increase mean field causes $\alpha_{EE}$ to become steeper, but $\alpha_{BB}$ steepens more dramatically.
@@ 4 Conclusions
In this work, we examine the $E-$mode and $B-$mode spectra from a suite of idealized, magnetized, and turbulent simulations. We find that isothermal turbulence alone is enough to reproduce the observed values of $\alpha_{EE}$ and $\alpha_{BB}$, as well as the ratio of amplitudes, $A_{BB}/A_{EE}$, for suitable values of Mach number, $\mathcal{M_{S}}$, and Alfvén Mach number, $\mathcal{M_{\rm{A}}}$. We additionally find that the observed correlation of $T$ and $E$, $r^{TE}=0.3$, is naturally reproduced by the turbulence at high $\mathcal{M_{S}}$ and an appropriate magnetic field strength. Correlations with $B$ are spectrally flat and near zero, certainly below $0.05$, but the results are somewhat murky. We suggest that a “typical” patch of the sky has $\mathcal{M_{S}}=4.7$, $\mathcal{M_{\rm{A}}}=1.5$, based on linear interpolation of the $E$- and $B$-mode slopes, but $r^{\rm{TE}}$ prefers lower $\mathcal{M_{\rm{A}}}$ and higher $\mathcal{M_{S}}$.
The density spectrum is found to be tightly related to $\mathcal{M_{S}}$, with slope $\alpha_{\rho}\sim-3.6+0.16\mathcal{M_{S}}$. This is due to the fact that shock thickness decreases with $\mathcal{M_{S}}$, leading to smaller scale structure and faster growth of instabilities such as Richmeyer-Meshkov and Rayleigh Taylor. The velocity spectra is relatively insensitive to $\mathcal{M_{S}}$ for $\mathcal{M_{S}}<2$. Supersonic slope values cluster around $\alpha_{v}\sim-3.5$ for super-Alfvén values, slightly shallower than the Kolmogorov value of -11/3. Magnetic spectral slopes are relatively insensitive to $\mathcal{M_{S}}$, and decrease with decreasing magnetic field.
@@
In Huffenberger et al. (2020), we model $E$ and $B$ with filaments that have an aspect ratio $\epsilon\in[0,1]$ threaded by magnetic fields at an angle, $\theta_{LH}$. It is found that as $\epsilon$ increases, the ratio of $B$ to $E$ increases as a longer filaments have proportionally more $E$. This also explains the decrease in $r^{TE}$ as $\epsilon$ increases. These predictions from the filament model are consistent with our findings with the turbulent boxes. As $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ increase, the ability of the magnetic field to suppress instability decreases, and shorter filaments are expected. This can be seen in projections and in the power spectra of $\rho.$ Commensurate, we find an increase in $B/E$ and a decrease in $r^{TE}$ with increasing $\mathcal{M_{S}}$. Again, the shorter structures as a result of increased $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ are in agreement, at least qualitatively, with the model of Huffenberger et al. (2020). Clark et al. (2021) model the parity violating correlation as a misalignment between filamentary structure and magnetic field direction in a similar filamentary framework, and compare to simulations. In future, we will examine filamentary properties of these cubes to further explore the predictive power of Huffenberger et al. (2020) and Clark et al. (2021).
In future simulation studies, we need higher resolution to increase our inertial range and improve our accuracy, particularly on the slopes. We also need larger statistical ensembles to quantify the sample variance in the $TB$ and $EB$ correlations. These are important for measurements of detector calibration, gravitational lensing, and cosmic birefringence.
@@ Acknowledgements
Support for this work was provided in part by the National Science Foundation under Grants AST-1616026 and AST-2009870 and NASA under 16-ATP16-0132. Simulations were performed on Stampede2, part of the Extreme Science and Engineering Discovery Environment (XSEDE; Towns et al., 2014), which is supported by National Science Foundation grant number ACI-1548562, under XSEDE allocation TG-AST140008.

### v2

@@ 3.2 Projected Intensity and Polarization Power Spectra
The $T$ spectrum, $C_{k}^{TT}$, can be seen in the first panel of Figures 6 and its slope, $\alpha_{TT}$, in the first panel of Figure 7. The $T$ spectrum has a slope that is nearly identical to the $\rho$ spectrum (though they appear different due to the difference in compensation). This is expected as in this model, $T$ is simply the projection of $\rho$. By the slice-projection theorem, a quantity $q(x,y,z)$, its projection $Q(x,y)$, and their Fourier transforms $\hat{q}(k_{x},k_{y},k_{z})$ and $\hat{Q}(k_{x},k_{y})$ are related as
@@
Interpreting $\alpha_{EE}$ (and $\alpha_{BB}$, next section) is tricky, since unlike the previous quantities, we lack even an adjacent theory about its behavior in a turbulent medium. It set by a combination of the geometry of density and magnetic structures, convolved with a kernel (Rotti & Huffenberger, 2019). We revisit this interpretation in Section 4.
@@
Figure 9: (Top) Correlation ratio $r_{xy}$ for TE, TB, and EB spectra. Vertical bars show the analysis window. (Bottom) Average $r_{XY}$ in the analysis window versus sonic Mach number, $\mathcal{M_{S}}$, compared to Planck-measured values. In the top plots, we use dotted lines for $0.70.4<\mathcal{M_{\rm{A}}}<$ 0.7 (smallest point size at bottom), dashed lines for $1.20.7<\mathcal{M_{\rm{A}}}<$ 1.2, and solid lines for $2.21.2<\mathcal{M_{\rm{A}}}<$ 2.2 (largest point size at bottom). See text for cautions about interpreting the $\langle r_{TB}\rangle$ and $\langle r_{TE}\rangle$.
@@ 3.3 Power Amplitude Ratios
The ratio of $B$ power to $E$ power is interesting because the observed value of $A_{\rm{BB}}/A_{\rm{EE}}=0.53$ was unexpected. We can also examine the ratio of of $E$ and $B$ relative to total power, $A_{\rm{EE}}/A_{\rm{TT}}$ and $A_{\rm{BB}}/A_{\rm{TT}}$, though our model lacks a detailed treatment of the dust polarization fraction. This implies that these ratios cannot be directly compared to observed values, though the trend with $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ can be measured. The unknown polarization fraction cancels out in the $A_{\rm{BB}}/A_{\rm{EE}}$ ratio, so this can be directly compared to the Planck value.
Figure 8 shows the ratio of fit amplitudes, $A_{\rm{EE}}/A_{\rm{TT}}$, $A_{\rm{BB}}/A_{\rm{TT}}$, and $A_{\rm{BB}}/A_{\rm{EE}}$ versus the sonic Mach number in each of the three panels. The third panel also shows the observed value as a grey line. The runs with slower velocities show the most $E$ modes, as the low velocity causes the field to have larger impact on the morphology, and more filamentary structures that align magnetic field and density are observed. This can be seen in the projections in Figure 1. For higher velocity ($\mathcal{M_{S}}\geq 4$), the amplitude ratios depend little on either the fluid velocity or the magnetic field. For $\mathcal{M_{S}}>4$, the ratio tends toward $A^{BB}/A^{EE}=0.55\pm 0.07$, near to the observed ratio of Planck. Thus it may be that that the ratio of $B$ to $E$ observed by Planck is a natural consequence of compressive turbulence, and that $A_{\rm{BB}}/A_{\rm{EE}}=0.53$ simply because the flow is hypersonic and magnetized, which naturally gives this value.
The study by Caldwell et al. (2017) used a linearization of the MHD equations to predict the amplitude ratio but had trouble reproducing the observed value. They postulate that the reason is the lack of nonlinearity in their treatment. Here we tend to agree, as our simulations become more nonlinear the observed slope is recovered.
@@ 3.4 Cross-correlations
for all pairs of $T$, $E$, and $B$. The top row shows the spectra, with the fitting window denoted by vertical light-grey lines. The bottom row shows $\langle r^{\rm{XY}}\rangle$, averaged over the fit window and all frames, as a function of sonic Mach number. Error bars are found by first averaging $r_{k}^{XY}$ over $k$ for each frame within the fit window, then taking the variance over frames. Shown in horizontal grey lines are the observed values of $r^{\rm{TE}}=0.355$ and $r^{\rm{TB}}=0.055$ (Planck Collaboration et al., 2020).
For the two even-parity modes, $T$ and $E$, we find significant positive correlation in all cases but one. As the velocity decreases and magnetic field increases, $r^{\rm{TE}}$ increases well above the value observed by Planck. The general trend is of increasing correlation with increasing field strength. For large sonic Mach numbers, correlations are more modest and mostly consistent with the observed Planck value of $r^{\rm{TE}}=0.35$. For large $\mathcal{M_{S}}$, the effect of the magnetic field is not as pronounced as it is for lower $\mathcal{M_{S}}$. Earlier results, e.g. for $\alpha_{EE}$ and the ratio of $E$ to $B$, show that the Planck data are consistent with $\mathcal{M_{S}}>4$, and this is also compatible with what we see for $\langle r^{\rm{TE}}\rangle$. We do observe a scale dependence in $r^{\rm{TE}}$ (rising toward small scales) that is not seen by Planck, which may be a sign that our simplified simulations are missing a key element of the ISM.
How do we draw conclusions from the $\langle r^{\rm{TE}}\rangle$ values, compared to the Planck value of 0.35? The Planck correlation looks most compatible with slow velocities with moderate magnetic fields or fast velocities with moderate-to-strong magnetic fields. Earlier we saw that the $B/E$ power ratio prefers fast velocities with $\mathcal{M_{S}}>4$, so the latter case seems to fit the bill. Other combinations do not work: in our simulations, low $\mathcal{M_{S}}$ and low $\mathcal{M_{\rm{A}}}$ (low velocity and strong magnetic field) would lead to a stronger correlation than what Planck sees. Low $\mathcal{M_{S}}$ and high $\mathcal{M_{\rm{A}}}$ (low velocity and weak magnetic field) leads to too small of a correlation, or a slight anticorrelation. We note that these simulations are idealized, and more realistic simulations may contain physical effects that reduce the filament alignment to the magnetic field and thus see the $TE$ correlations reduced.
@@
These tendencies toward positive parity-violating correlation values at high fluid velocity are somewhat puzzling because all of the MHD physics we include respects parity. The simulations start with uniform density and are driven with a non-helical acceleration pattern, and the base solver is an unsplit solver with no inherent asymmetry. A larger, more systematic, and more resolved ensemble of simulations will be necessary to determine if we have inadvertently inserted some parity-violating effect, if this is simply a statistical fluctuation, or if there is a slight tendency for such MHD simulations to produce positive parity violations.
@@ 3.5 Linear fits and importance of parameters
where $q$ stands for density, velocity, magnetic field, $T$, $E$, and $B$. These fit coefficients are found in Table 1. Noting that $b=\partial q/\partial\mathcal{M_{S}}$ and $c=\partial q/\partial\mathcal{M_{\rm{A}}}$, we see that $b$ and $c$ give the relative importance of sonic and Alfvén Mach numbers on each quantity. The third column of Table 1 gives the ratio of $c/b$, which denotes the relative impact of the two. Density, $\rho$, depends only on sonic Mach number $(b_{\rho}=0)$ while the velocity spectrum is more influenced by $\mathcal{M_{\rm{A}}}$ ($c_{\rho}/b_{\rho}=6.5$). Sonic Mach number determines $\alpha_{TT}$, while Alfvén Mach number determines $\alpha_{BB}$.
@@ 3.6 Parallel versus perpendicular projections
Figure 11 shows $\alpha_{EE}$ and $\alpha_{BB}$ vs $\mathcal{M_{S}}$ for parallel projections (top row) and perpendicular projections (bottom row, same as Figure 7, reproduced for ease of comparison). For the weakly magnetized cases (large points) the behavior is comparable between the two directions, as expected. For the more strongly magnetized case, increasing magnetic field has the opposite effect on the slope between the two directions. For the $\hat{x}$ projection, increasing magnetic field makes $\alpha_{EE}$ and $\alpha_{BB}$ slightly more shallow. For the parallel direction, increase mean field causes $\alpha_{EE}$ to become steeper, but $\alpha_{BB}$ steepens more dramatically.
@@
4 Discussion
Here we discuss our results in the context of our model foregrounds as an ensemble of filaments. In Huffenberger et al. (2020), we model $E$ and $B$ with filaments that have an aspect ratio $\epsilon\in[0,1]$ threaded by magnetic fields at an angle, $\theta_{LH}$. As $\epsilon$ increases, making the filaments more round, the ratio of $B$ to $E$ increases as a shorter, rounder filaments have proportionally less $E$ (see Figure 6 of Huffenberger et al. 2020). This also explains the decrease in $r^{TE}$ as $\epsilon$ increases. These predictions from the filament model are consistent with our findings with the turbulent boxes. As $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ increase, the ability of the magnetic field to suppress instability decreases, and shorter filaments are expected. This can be seen in projections and in the power spectra of $\rho.$ Commensurate, we find an increase in $B/E$ and a decrease in $r^{TE}$ with increasing $\mathcal{M_{S}}$. Again, the shorter structures as a result of increased $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ have effects on the polarization power spectra that are in agreement, at least qualitatively, with the model of Huffenberger et al. (2020). Clark et al. (2021) model the parity violating correlation as a misalignment between filamentary structure and magnetic field direction in a similar filamentary framework, and compare to simulations. In future, we will examine filamentary properties of these cubes to further explore the predictive power of Huffenberger et al. (2020) and Clark et al. (2021).
@@ 5 Conclusions
In this work, we examine the $E-$mode and $B-$mode spectra from a suite of idealized, magnetized, and turbulent simulations. We find that isothermal turbulence alone is enough to reproduce the observed values of $\alpha_{EE}$ and $\alpha_{BB}$, as well as the ratio of amplitudes, $A_{BB}/A_{EE}$, for suitable values of sonic Mach number, $\mathcal{M_{S}}$, and Alfvén Mach number, $\mathcal{M_{\rm{A}}}$. We additionally find that the observed correlation of $T$ and $E$, $r^{TE}=0.3$, is naturally reproduced by the turbulence at high $\mathcal{M_{S}}$ and an appropriate magnetic field strength. Parity-violating correlations with $B$ are spectrally flat and near zero, certainly below $0.05$, but the results are somewhat murky. We suggest that a “typical” patch of the sky has $\mathcal{M_{S}}=4.7$, $\mathcal{M_{\rm{A}}}=1.5$, based on linear interpolation of the $E$- and $B$-mode slopes, but $r^{\rm{TE}}$ prefers lower $\mathcal{M_{\rm{A}}}$ and higher $\mathcal{M_{S}}$.
The density spectrum is found to be tightly related to $\mathcal{M_{S}}$, with slope $\alpha_{\rho}\sim-3.6+0.16\mathcal{M_{S}}$. This is due to the fact that shock thickness decreases with $\mathcal{M_{S}}$, leading to smaller scale structure and faster growth of instabilities such as Richmeyer-Meshkov and Rayleigh Taylor. The velocity spectra is relatively insensitive to $\mathcal{M_{S}}$ for $\mathcal{M_{S}}>2$. Supersonic slope values cluster around $\alpha_{v}\sim-3.5$ for super-Alfvén values, slightly shallower than the Kolmogorov value of $-11/3$. Magnetic spectral slopes are relatively insensitive to $\mathcal{M_{S}}$, and decrease with decreasing magnetic field.
@@
In future simulation studies, we need higher resolution to increase our inertial range and improve our accuracy, particularly on the slopes. We also need larger statistical ensembles to quantify the sample variance in the $TB$ and $EB$ correlations. These parity-violating correlations are important for measurements of detector calibration, gravitational lensing, and cosmic birefringence.
@@ Acknowledgements
Support for this work was provided in part by the National Science Foundation under grants AST-1616026 and AST-2009870, NASA under grant NNX17AF87G and 80NSSC23K0466, and the DOE under grant DE-SC0024462. Simulations were performed on Stampede2, part of the Extreme Science and Engineering Discovery Environment (XSEDE; Towns et al., 2014), which is supported by National Science Foundation grant number ACI-1548562, under XSEDE allocation TG-AST140008. We thank the anonymous referee for several suggestions to clarify the text.